OpenStudy (anonymous):

rewrite ln l sin theta l -ln l cos theta l as a single logarithm and then simplify the result

OpenStudy (anonymous):

ln(a) - ln(b) = ln(a/b) sin(x)/cos(x)=tan(x)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Have anymore questions?

OpenStudy (anonymous):

only other problems like partial fraction decomp

OpenStudy (anonymous):

I can help if you'd like

OpenStudy (anonymous):

ok thanks find the partial fraction decomp of (5x+3)/(x^2-3x-10)